OpenStudy (anonymous):

simplify and reduce: 3 3/10 + 1 1/10

OpenStudy (anonymous):

Is this your question?\[\frac{ 33 }{ 10 } + \frac {11}{10}\]

OpenStudy (ajprincess):

or this \[3\frac{ 3 }{ 10}+1\frac{ 1 }{ 10 }?\]

OpenStudy (anonymous):

the second one ^^

OpenStudy (anonymous):

4 2/5

OpenStudy (anonymous):

thanks!